OpenStudy (anonymous):

Derive this series expansion: (I was wondering if there was a way to do it other than the long way) ln((1+x)/(1-x)) = 2(x + (x^3/3) + (x^5/5) + ...)

OpenStudy (anonymous):

i would start with \[\ln(1+x)-\ln(1-x)\]

OpenStudy (anonymous):

That makes sense.

OpenStudy (anonymous):

the expansion for \[ln(1+x)\] is well known as \[x-\frac{x^2}{2}+\frac{x^3}{3}-\frac{x^4}{4}+...\] then you can replace x by -x to get the expansion for \[\ln(1-x)\] and then add term by term

OpenStudy (anonymous):

yep. thanks.

OpenStudy (anonymous):

yw